can you help by finding each quotient:
6 1/4 divided by 2 1/2=
2 1/3 divided by 7/10=
Also can you explain me the process to get the proper answer?
Thank you
Isaias

Answer by gonzo(654)

problem 1:
6 and 1/4 divided by 2 and 1/2.
6 and 1/4 = 25/4
2 and 1/2 = 5/2
the problem becomes 25/4 divided by 5/2.
25/4 divided by 5/2 is the same as 25/4 * 2/5.
you take the inverse of the divisor and make it a multiplier.
if you're dividing by 5, you would then multiply by 1/5.
if you're dividing by 1/5, you would then multiply by 5.
so, 25/4 divided by 5/2 is the same as 25/4 multiplied by 2/5.
25/4 multiplied by 2/5 = 25*2/4*5 = 50/20 which is your answer to the first problem. you can simplify it by removing common factors in the numerator and the denominator so the answer would be 5/2 in its reduced form.
problem 2:
same deal, different numbers.
2 and 1/3 divided by 7/10.
2 and 1/3 = 7/3.
7/10 = 7/10 (stays the same).
7/3 divided by 7/10 is the same as 7/3 multiplied by 10/7.
7/3 multiplied by 10/7 = 70/21 which is your answer to the second problem. you can also reduce this to simplest terms by eliminating the common factors in the numerator and the divisor. the number then reduces to 10/3.
you can check your answers as follows:
problem 1:
answer was 5/2.
multiply your answer by the number you were originally dividing by.
5/2 multiplied by 5/2 = 25/4.
25/4 is the same as 6 and 1/4 which was the original numerator you started out with.
problem 2:
answer was 10/3.
multiply your answer by the number you were originally dividing by.
10/3 multiplied by 7/10 = 70/30 = 7/3 = 2 and 1/3 which is the numerator you originally started with.